A solution is composed of one mole of non-volatile solute and four moles of volatile solvent. If the vapor pressure of the pure solvent is 20 torr, what is the vapor pressure of the solution?
1
Expert's answer
2014-05-15T13:14:05-0400
The solute is non-volatile, so the vapor pressure will be equal to the pressure of the solvent.
According to Raoult's law:
Pi = xiPoi
where xi is a mole fraction, Poi is the pressure of pure substance.
Calculate the mole fraction of the solvent:
x(solvent) = n(solvent)/[n(solvent) + n(solute)] = 4/5 = 0.8
P(solvent) = 0.8*20 = 16 torr
Answer: P = 16 torr.